OpenStudy (anonymous):

simplify (2 √5+3 √7)^2 show your work

OpenStudy (anonymous):

\[(2\sqrt{5} + 3\sqrt{7})^{2} = 20 + 12\sqrt{35} + 63 = 83 + 12\sqrt{35}\]

OpenStudy (anonymous):

You basically foil it.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome.